OpenStudy (anonymous):

find the period and amplitude of the periodic function. y = -3 cos 4x

Nnesha (nnesha):

\[\huge\rm y = -3 \cos 4x\]

Nnesha (nnesha):

\[\huge\rm y = A \cos (B \theta -C) + D \] this is equation where A is amplitude and to find period you have to divide 2pi/B

Directrix (directrix):

@fujiblu According to the formula, the period is 2π/ 4 = ? when simplified.

Directrix (directrix):

And, amplitude is | -3 | = ?